### v2

@@ 5.2 Comparison with Gurobi, dADMM and ES-iALM
1000 3.89e-07 32 4.63 3.21e-05 10000 16.11 8.33e-05 16 (15000) 18.87 6.44e-07 18 (67) 1.93
2000 2.71e-07 36 21.88 1.26e-04 10000 102.04 7.87e-05 15 (15000) 107.03 6.49e-07 18 (77) 9.69
3000 3.88e-07 40 57.76 3.10e-04 10000 243.63 8.38e-05 15 (15000) 255.45 6.73e-07 19 (86) 27.50
4000 3.81e-07 41 120.33 5.90e-04 10000 439.32 6.54e-05 14 (15000) 461.86 6.81e-07 19 (93) 56.32
5000 5.93e-07 47 201.37 7.22e-04 10000 689.08 3.25e-05 15 (15000) 729.19 6.18e-07 19 (99) 97.73
6000 5.47e-07 48 308.46 1.12e-03 10000 994.07 3.19e-04 14 (15000) 1047.92 6.50e-07 20 (110) 155.45
7000 4.46e-07 46 409.37 1.40e-03 10000 1354.34 1.67e-04 14 (15000) 1420.33 6.10e-07 20 (116) 225.95
8000 1.37e-06 47 615.47 1.80e-03 10000 1762.95 3.67e-04 14 (15000) 1852.26 6.25e-07 20 (120) 306.10
9000 8.38e-07 50 779.09 2.04e-03 10000 2230.05 1.84e-04 14 (15000) 2343.77 6.82e-07 20 (127) 408.59
10000 7.11e-07 52 1051.46 1.99e-03 10000 2750.64 1.44e-04 14 (15000) 2897.32 7.40e-07 20 (128) 504.35
Figure 1: Comparisons among Gurobi, dADMM, ES-iALM, and ripALM for solving the QROT problem (4.1) with $m=n\in\{1000,\,2000,\cdots,10000\}$. The shaded region indicates the maximum and minimum computational times taken to solve 10 instances for each problem dimension.
5.3 Experiments for BPDN
@@
In this section, we compare our ripALM with Mosek (version 11.0.27), SPGL1 (version 2.1), and dADMM (see Appendix E) for solving the BPDN problem (4.9). For ripALM, we set $\rho=0.99$ as in the previous section. For dADMM, we initialize the penalty parameter as $\sigma=1$, and dynamically update it based on the primal-dual residuals, following the strategy in [57, Section 5.1]. We terminate dADMM when it returns a point $(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})$ satisfying $\Delta_{\mathrm{res}}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})<10^{-6}$ or reaches 10000 iterations. For SPGL1, we use its default termination conditions, set the tolerance as $10^{-6}$, (aligning it with our tolerance), and set the maximum number of iterations as 10000. For Mosek, we use its default termination conditions, but adopt two tolerance levels: a tolerance of $10^{-6}$ (denoted as Mosek (1e-6)) to match our tolerance and a tighter tolerance of $10^{-10}$ (denoted as Mosek (1e-10)) to generate a high-precision solution to use as a benchmark.
We construct the tested BPDN problems using both synthetic datasets and UCI datasets [^6] Available at: https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/regression.html. The synthetic datasets are generated as follows. Given a triple $(m,n,s)$, we first generate a dictionary matrix $D\in\mathbb{R}^{m\times n}$ with i.i.d. standard Gaussian entries. We then uniformly select an index set $\mathcal{S}\subset\{1,\cdots,n\}$ with cardinality $|\mathcal{S}|=s$ and construct an $s$-sparse signal $\bar{\bm{s}}\in\mathbb{R}^{n}$, whose entries on $\mathcal{S}$ are drawn independently from the standard Gaussian distribution and are zero elsewhere. The measurement vector is then formed as $\bm{b}=D\bar{\bm{s}}+\delta\bm{\zeta}$, where $\bm{\zeta}\in\mathbb{R}^{m}$ is a noise vector with i.i.d. standard Gaussian entries and $\delta>0$ is the noise level. For UCI datasets, we choose housing7 and mpg7 to obtain $(D,\bm{b})$, where $D$ has dimensions $506\times 77520$ for housing7 and $392\times 3432$ for mpg7. Note that housing7 and mpg7 are expanded versions of their datasets, with the suffix ‘7’ indicating that an order $7$ polynomial is used to generate the basis functions; see [22, Section 4.1] for more details. Finally, to evaluate the solution quality, we compute the relative feasibility residual and normalized objective gap, defined respectively as:
@@
$\texttt{feas}(\bm{s}):=\frac{\max\left\{\|D\bm{s}-\bm{b}\|_{2}-\widehat{\kappa},\,0\right\}}{1+\|\bm{b}\|_{2}},\quad\text{ and }\quad\texttt{nobj}(\bm{s}):=\frac{\big|\|\bm{s}\|_{1}-\|\bm{s}^{*}\|_{1}\big|}{1+\|\bm{s}^{*}\|_{1}},$
@@
Table 5 reports the average computational results for (m,n,s)=(0.1i,i, 0.02i)(m,n,s)=(0.1i,\,i,\,0.02i) with $i\in\{10000,20000,30000\}$ over $10$ randomly generated BPDN instances. In these experiments, the parameter $\widehat{\kappa}$ in (4.9) is set to $\widehat{\kappa}=\delta\left\|\bm{\zeta}\right\|_{2}$ with $\delta\in\{0.01,\,0.1\}$. The results demonstrate the superior performance of ripALM: it consistently achieves the smallest normalized objective gap, while requiring the least computational time and maintaining favorable feasibility residuals. Table 6 reports the results on two UCI data sets, where we choose $\widehat{\kappa}=\delta\|\bm{b}\|$ with $\delta\in\{0.01,\,0.1\}$. These results further show the encouraging performance of our ripALM on the real-world datasets.
Table 5: Numerical results of Mosek, SPGL1, dADMM, and ripALM for solving the BPDN problem (4.9) with $(m,n,s)=(0.1i,i,0.02i)$ and $\widehat{\kappa}=\delta\left\|\bm{\zeta}\right\|_{2}$. In the table, “nobj” denotes the normalized objective gap; “feas” denotes the relative feasibility residual; “#” denotes the number of iterations (the total number of Ssn iterations is given in the bracket); “time” denotes the computational time.
$\delta=0.01$ $\delta=0.1$
$i$ method nobj feas # time nobj feas # time
10000 Mosek (1e-10) 0 7.94e-12 26 6.80 0 1.04e-11 24 6.31
Mosek (1e-6) 1.09e-05 3.42e-07 19 5.52 1.00e-05 3.12e-07 18 5.21
SPGL1 4.72e-05 0 4520 28.32 1.35e-05 0 703 6.15
dADMM 7.49e-06 4.49e-12 10000 71.65 1.95e-06 8.98e-13 4081 31.20
ripALM 1.35e-08 1.31e-09 9 (58) 2.26 4.35e-09 2.89e-09 9 (46) 1.87
20000 Mosek (1e-10) 0 7.50e-12 32 36.20 0 8.51e-12 28 31.61
Mosek (1e-6) 1.12e-05 3.42e-07 21 25.79 1.05e-05 3.20e-07 20 25.02
SPGL1 4.83e-05 0 7991 219.83 3.71e-05 0 997 34.68
dADMM 8.75e-06 1.83e-12 10000 286.40 1.93e-06 5.38e-13 5391 179.74
ripALM 2.83e-09 7.64e-10 9 (81) 12.63 1.06e-08 3.47e-10 10 (62) 10.17
30000 Mosek (1e-10) 0 9.16e-12 30 83.94 0 7.44e-12 32 87.60
Mosek (1e-6) 8.35e-06 2.51e-07 22 67.40 1.31e-05 3.95e-07 21 64.39
SPGL1 5.28e-05 0 8156 479.54 1.69e-05 0 1038 80.32
dADMM 9.36e-06 2.06e-12 10000 601.85 1.88e-06 2.64e-13 6121 453.56
ripALM 5.28e-09 7.82e-10 8 (81) 32.52 4.49e-08 3.88e-10 9 (71) 28.50
Table 6: Same as Table 5 but for the UCI datasets with $\widehat{\kappa}=\delta\|\bm{b}\|$.
$\delta=0.01$ $\delta=0.1$
datasets method nobj feas # time nobj feas # time
housing7 Mosek (1e-10) 0 2.37e-11 36 44.94 0 2.40e-12 27 35.05
Mosek (1e-6) 1.83e-05 4.60e-08 22 25.42 4.31e-05 1.44e-08 19 23.89
SPGL1 1.76e-01 1.42e-02 10000 254.94 9.21e-03 0 1539 51.72
dADMM 4.98e-06 0 10000 237.17 1.74e-06 0 3201 76.53
ripALM 6.58e-06 0 22 (74) 2.28 2.08e-09 0 18 (45) 1.15
mpg7 Mosek (1e-10) 0 1.49e-09 26 1.03 0 1.52e-13 16 0.65
Mosek (1e-6) 1.45e-07 2.98e-08 17 0.62 8.92e-06 1.82e-08 13 0.53
SPGL1 8.82e-01 3.85e-02 10000 3.00 5.35e-03 0 443 0.22
dADMM 1.48e-04 0 10000 3.05 2.33e-06 0 2001 0.60
ripALM 1.03e-06 1.73e-08 30 (126) 0.49 2.06e-07 1.58e-10 18 (23) 0.02
@@ 6 Conclusions
In this paper, we developed a relative-type inexact proximal augmented Lagrangian method (ripALM) for solving a class of linearly constrained convex optimization problems. The proposed ripALM is the first relative-type inexact version of the vanilla pALM with provable convergence guarantees. By employing a suitable relative-type error criterion, it simplifies implementation and parameter tuning, compared to the absolute-type inexact pALM. We conducted a thorough convergence analysis and demonstrated the competitive efficiency of our ripALM through numerical experiments on solving quadratically regularized optimal transport problems and basis pursuit denoising problems. Since our model covers a wide range of applications, future work may explore the potential of applying ripALM to other practical problems, including general regularized optimal transport problems as studied in [58]. Additionally, it would be interesting to extend ripALM to a broader class of convex optimization problems, thereby further enhancing its applicability.
@@ Acknowledgements
We thank the associate editor and two anonymous reviewers for their valuable comments and suggestions, which have helped to improve the quality of this paper. The research of Lei Yang is supported in part by the National Key Research and Development Program of China under grant 2023YFB3001704 and the National Natural Science Foundation of China under grant 12301411. The research of Kim-Chuan Toh is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 grant (MOE-T2EP20224-0017).
@@ Appendix A Missing Proofs in Section 3
Since $\{\nu_{k}\}$ is a non-negative summable sequence, it then follows from [35, Lemma 2 in Section 2.2 on Page 44] that the sequence $\left\{\|\bm{x}^{k}-\bm{x}^{*}\|^{2}+\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}\right\}$ is convergent (but do not necessarily converge to 0). This together with $\tau_{k}\geq\tau_{\min}>0$ implies that all sequences $\{\bm{x}^{k}\}$, $\{\bm{w}^{k}\}$, $\{\bm{y}^{k}\}$ are bounded.
@@
Note that $F$ is closed upper semicontinuous concave (see, for example, [40, Theorem 7]), and thus $\mathrm{dom}\,F$ is closed. This together with $(\bm{0},\,\bm{x}^{\infty})=\lim\limits_{i\to\infty}\big(\bm{\theta}^{k_{i}+1},\,\bm{x}^{k_{i}+1}\big)$ induces that $(\bm{0},\,\bm{x}^{\infty})\in\mathrm{dom}\,F$. Then, we see that
@@
We next study the limit of $\{G(\bm{y}^{k+1},\bm{\xi}^{k+1})\}$. Since F-F and $G$ are convex conjugate and $\big(\bm{\theta}^{k+1},\,\bm{x}^{k+1}\big)\in\partial G(\bm{y}^{k+1},\,\bm{\xi}^{k+1})$, we can get the following equality by using the Fenchel equality (see, for example, [39, Theorem 23.5]):
@@
Then, by passing to a further subsequence if necessary, we may also assume without loss of generality that the subsequence $\{\bm{x}^{k_{j}}\}\subseteq\{\bm{x}^{k}\}$ converges to some accumulation point $\bm{x}^{\infty}$, which, in view of statement (iv), belongs to $\mathcal{X}^{*}$ (the optimal solution set of (3.2)). Thus, for such $\bm{x}^{\infty}$ and any $\bm{y}^{*}\in\mathcal{Y}^{*}$, using (A.10) with some manipulations, we can obtain that, for all $k>k_{j}$,
@@
Since $0\leq\phi\leq\liminf\limits_{k\to\infty}\left\{\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}\right\}$ for any $\bm{y}^{*}\in\mathcal{Y}^{*}$, passing to the limit superior when $k\to\infty$ on the both sides of the above inequality, we obtain that, for any $\bm{y}^{*}\in\mathcal{Y}^{*}$,
@@
where $\mu_{k}:=\sqrt{\frac{1+\nu_{k}}{1+\gamma_{k}}}$. Since $\nu_{k}\to 0$ and $\gamma_{k}\geq\gamma_{\min}>0$ for all $k\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\geq k_{0}$, one can verify that $\limsup\limits_{k\to\infty}\,\{\mu_{k}\}<1$. Thus, we obtain the desired results in statement (i).
@@
$\widetilde{c}\,\tau_{\min}\|\bm{y}^{k+1}-{\bm{y}}^{k}\|^{2}\leq\widetilde{c}\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-{\bm{y}}^{k})\\ \bm{x}^{k+1}-{\bm{x}}^{k}\end{matrix}\right\|^{2}\leq\mathrm{dist}_{\Lambda^{k}}^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}2}\left((\bm{y}^{k},\bm{x}^{k}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right),$

### v3

@@ 5.1 Experiments on QROT
1000 4.74e-07 32 4.98 3.17e-05 10000 17.74 4.74e-05 16 (15000) 17.77 6.37e-07 18 (66) 2.12
2000 3.61e-07 38 22.00 1.39e-04 10000 106.15 1.55e-05 15 (15000) 105.57 5.96e-07 18 (78) 11.08
3000 4.69e-07 40 56.67 2.78e-04 10000 251.36 1.72e-05 15 (15000) 254.22 6.41e-07 19 (84) 29.38
4000 4.20e-07 43 114.19 5.01e-04 10000 450.94 4.11e-05 15 (15000) 454.35 6.37e-07 19 (93) 61.01
5000 5.39e-07 46 194.45 7.26e-04 10000 706.19 4.63e-05 15 (15000) 716.31 6.58e-07 19 (101) 109.55
6000 1.33e-06 45 278.79 1.08e-03 10000 1017.11 6.80e-05 15 (15000) 1030.13 6.47e-07 20 (109) 171.26
7000 8.21e-07 47 395.60 1.30e-03 10000 1382.91 1.08e-04 14 (15000) 1399.50 5.65e-07 20 (117) 255.10
8000 7.15e-07 48 604.52 1.62e-03 10000 1803.87 1.53e-04 14 (15000) 1824.82 6.29e-07 20 (119) 333.22
9000 6.89e-07 50 747.40 2.15e-03 10000 2283.73 2.01e-04 14 (15000) 2309.05 5.80e-07 21 (129) 472.94
10000 1.10e-06 52 1026.56 2.43e-03 10000 2814.83 2.23e-04 14 (15000) 2849.56 5.01e-07 21 (137) 626.74
5.2 Experiments on BPDN
In this section, we evaluate the performance of different algorithms for solving the BPDN problem (4.9). We use the datasets housing7 and mpg7 from the UCI data repository [^5] Available at: https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/regression.html to generate the test instances $(D,\bm{b})$, where $D$ has dimensions $506\times 77520$ for housing7 and $392\times 3432$ for mpg7. Here, housing7 and mpg7 are expanded versions of the original datasets, with the suffix ‘7’ indicating that a polynomial of degree 7 is used to generate the basis functions; see [17, Section 4.1] for more details. The parameter $\widehat{\kappa}$ in (4.9) is set to $\widehat{\kappa}=\delta\|\bm{b}\|$ with $\delta\in\{0.01,\,0.1\}$.
We first examine how different error criteria and tolerance parameters influence the numerical performance of different inexact pALMs. The tolerance parameters are chosen in the same way as in Table 2. Table 4 presents the results for housing7. As in the QROT experiments, the performance of all three algorithms is affected by the choice of tolerance parameters, while ripALM and ciPALM require less tuning effort as they involve only a single tolerance parameter $\rho\in[0,1)$. Moreover, ripALM consistently outperforms ciPALM in terms of the total number of Ssn iterations and shows better efficiency and robustness.
Table 4: Comparisons among Snipal, ciPALM and ripALM under different choices of tolerance parameters for solving BPDN problem (4.9) with $\widehat{\kappa}=\delta\|\bm{b}\|$ and $\delta\in\{0.1,0.01\}$, using housing7 from the UCI data repository.
$\delta=0.1$
Snipal ciPALM ripALM
$(\varepsilon_{0}\!=\!\delta_{0},p,q)$ # (#) $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ # (#) $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ # (#) $\rho$ # (#) $\rho$ # (#)
(1, 1.1, 1.1) 19 (71) (0.1, 1.1, 1.1) 19 (78) (0.01, 1.1, 1.1) 19 (84) 0.999 19 (149) 0.999 19 (67)
(1, 1.1, 2.1) 19 (75) (0.1, 1.1, 2.1) 19 (84) (0.01, 1.1, 2.1) 19 (86) 0.99 19 (149) 0.99 19 (67)
(1, 1.1, 3.1) 19 (78) (0.1, 1.1, 3.1) 19 (85) (0.01, 1.1, 3.1) 19 (94) 0.9 19 (149) 0.9 19 (64)
(1, 2.1, 1.1) 19 (75) (0.1, 2.1, 1.1) 19 (87) (0.01, 2.1, 1.1) 19 (85) 0.7 19 (132) 0.7 19 (64)
(1, 2.1, 2.1) 19 (77) (0.1, 2.1, 2.1) 19 (84) (0.01, 2.1, 2.1) 19 (86) 0.5 19 (132) 0.5 19 (64)
(1, 2.1, 3.1) 19 (80) (0.1, 2.1, 3.1) 19 (90) (0.01, 2.1, 3.1) 19 (94) 0.3 19 (94) 0.3 19 (67)
(1, 3.1, 1.1) 19 (83) (0.1, 3.1, 1.1) 19 (84) (0.01, 3.1, 1.1) 19 (91) 0.1 19 (87) 0.1 19 (76)
(1, 3.1, 2.1) 19 (83) (0.1, 3.1, 2.1) 19 (84) (0.01, 3.1, 2.1) 19 (92) 0.05 18 (76) 0.05 19 (74)
(1, 3.1, 3.1) 19 (84) (0.1, 3.1, 3.1) 19 (90) (0.01, 3.1, 3.1) 19 (93) 0.01 19 (76) 0.01 19 (73)
$\delta=0.01$
Snipal ciPALM ripALM
$(\varepsilon_{0}\!=\!\delta_{0},p,q)$ # (#) $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ # (#) $(\varepsilon_{0}\!=\!\delta_{0},p,q)$ # (#) $\rho$ # (#) $\rho$ # (#)
(1, 1.1, 1.1) 24 (156) (0.1, 1.1, 1.1) 24 (159) (0.01, 1.1, 1.1) 24 (163) 0.999 24 (537) 0.999 24 (137)
(1, 1.1, 2.1) 24 (157) (0.1, 1.1, 2.1) 24 (161) (0.01, 1.1, 2.1) 24 (170) 0.99 24 (537) 0.99 24 (137)
(1, 1.1, 3.1) 24 (158) (0.1, 1.1, 3.1) 24 (165) (0.01, 1.1, 3.1) 24 (170) 0.9 24 (537) 0.9 24 (137)
(1, 2.1, 1.1) 24 (158) (0.1, 2.1, 1.1) 24 (170) (0.01, 2.1, 1.1) 24 (176) 0.7 24 (310) 0.7 24 (138)
(1, 2.1, 2.1) 24 (158) (0.1, 2.1, 2.1) 24 (170) (0.01, 2.1, 2.1) 24 (176) 0.5 24 (256) 0.5 24 (137)
(1, 2.1, 3.1) 24 (161) (0.1, 2.1, 3.1) 24 (172) (0.01, 2.1, 3.1) 24 (176) 0.3 24 (265) 0.3 24 (137)
(1, 3.1, 1.1) 24 (164) (0.1, 3.1, 1.1) 24 (176) (0.01, 3.1, 1.1) 24 (185) 0.1 24 (214) 0.1 24 (143)
(1, 3.1, 2.1) 24 (164) (0.1, 3.1, 2.1) 24 (176) (0.01, 3.1, 2.1) 24 (185) 0.05 24 (222) 0.05 24 (144)
(1, 3.1, 3.1) 24 (164) (0.1, 3.1, 3.1) 24 (176) (0.01, 3.1, 3.1) 24 (185) 0.01 24 (169) 0.01 24 (145)
@@ 5.2 Experiments on BPDN
We next compare ripALM with Mosek (version 11.0.27), SPGL1 (version 2.1), and dADMM (see Appendix E). For ripALM, we set $\rho=0.99$ based on results in Table 4. For dADMM, we initialize the penalty parameter as $\sigma=1$, and dynamically update it based on the primal-dual residuals, following the strategy in [47, Section 5.1]. We terminate dADMM when it returns a point $(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})$ satisfying $\Delta_{\mathrm{res}}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})<10^{-6}$ or reaches 10000 iterations. For SPGL1, we use its default termination conditions, set the tolerance as $10^{-6}$, (aligning it with our tolerance), and set the maximum number of iterations as 10000. For Mosek, we use its default termination conditions, but adopt two tolerance levels: a tolerance of $10^{-6}$ (denoted as Mosek (1e-6)) to match our tolerance and a tighter tolerance of $10^{-10}$ (denoted as Mosek (1e-10)) to generate a high-precision solution to use as a benchmark. In addition, to evaluate the solution quality, we compute the relative feasibility residual and normalized objective gap, defined respectively as:
@@
$\texttt{feas}(\bm{s}):=\frac{\max\left\{\|D\bm{s}-\bm{b}\|-\widehat{\kappa},\,0\right\}}{1+\|\bm{b}\|},\quad\text{ and }\quad\texttt{nobj}(\bm{s}):=\frac{\big|\|\bm{s}\|_{1}-\|\bm{s}^{*}\|_{1}\big|}{1+\|\bm{s}^{*}\|_{1}},$
@@
Table 5 reports the results for housing7 and mpg7. These results further demonstrate the effectiveness of ripALM in comparison with the other representative algorithms.
Table 5: Comparisons among Mosek, SPGL1, dADMM, and ripALM for solving the BPDN problem (4.9) with $\widehat{\kappa}=\delta\|\bm{b}\|$ and $\delta\in\{0.1,0.01\}$, using housing7 and mpg7 from the UCI data repository. In the table, “nobj” denotes the normalized objective gap; “feas” denotes the relative feasibility residual; “#” denotes the number of iterations (the total number of Ssn iterations is given in the bracket); “time” denotes the computational time.
nobj feas iters time nobj feas iters time
dataset method $\delta=0.1$ $\delta=0.01$
housing7 Mosek (1e-10) 0 2.40e-12 27 33.21 0 2.37e-11 36 43.57
Mosek (1e-6) 4.31e-05 1.44e-08 19 22.75 1.83e-05 4.60e-08 22 24.49
SPGL1 9.21e-03 0 1539 48.46 1.76e-01 1.42e-02 10000 237.72
dADMM 2.00e-06 0 3151 70.09 4.98e-06 0 10000 221.44
ripALM 2.08e-09 0 18 (48) 1.65 6.58e-06 0 22 (74) 2.11
mpg7 Mosek (1e-10) 0 1.52e-13 16 0.62 0 1.49e-09 26 0.99
Mosek (1e-6) 8.92e-06 1.82e-08 13 0.51 1.45e-07 2.98e-08 17 0.60
SPGL1 5.35e-03 0 443 0.21 8.82e-01 3.85e-02 10000 2.14
dADMM 2.33e-06 0 2001 0.70 1.48e-04 0 10000 3.24
ripALM 2.06e-07 1.58e-10 18 (23) 0.02 1.03e-06 1.73e-08 30 (126) 0.45
@@ 6 Conclusions
In this paper, we developed a relative-type inexact proximal augmented Lagrangian method (ripALM) for solving a class of linearly constrained convex optimization problems. To the best of our knowledge, the proposed ripALM is the first relative-type inexact version of the vanilla pALM with provable convergence guarantees. By employing a suitable relative-type error criterion, it simplifies implementation and parameter tuning, compared to the classical absolute-type inexact pALM. We conducted a thorough convergence analysis and demonstrated the competitive efficiency of ripALM through numerical experiments on solving quadratically regularized optimal transport problems and basis pursuit denoising problems. Since our model covers a wide range of applications, future work may explore the potential of applying ripALM to other practical problems.
@@ Acknowledgements
The research of Lei Yang is supported in part by the National Key Research and Development Program of China under grant 2023YFB3001704 and the National Natural Science Foundation of China under grant 12301411. The research of Kim-Chuan Toh is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 grant (MOE-T2EP20224-0017).
@@ Appendix A Missing Proofs in Section 3
Since $\{\nu_{k}\}$ is a non-negative summable sequence, it then follows from [30, Lemma 2 in Section 2.2 on Page 44] that the sequence $\left\{\|\bm{x}^{k}-\bm{x}^{*}\|^{2}+\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}\right\}$ is convergent (but do not necessarily converge to 0). This together with $\tau_{k}\geq\tau_{\min}>0$ implies that all sequences $\{\bm{x}^{k}\}$, $\{\bm{w}^{k}\}$, $\{\bm{y}^{k}\}$ are bounded.
@@
Note that $F$ is closed upper semicontinuous concave (see, for example, [36, Theorem 7]), and thus $\mathrm{dom}\,F$ is closed. This together with $(\bm{0},\,\bm{x}^{\infty})=\lim\limits_{i\to\infty}\big(\bm{\theta}^{k_{i}+1},\,\bm{x}^{k_{i}+1}\big)$ induces that $(\bm{0},\,\bm{x}^{\infty})\in\mathrm{dom}\,F$. Then, we see that
@@
We next study the limit of $\{G(\bm{y}^{k+1},\bm{\xi}^{k+1})\}$. Since F-F and $G$ are convex conjugate and $\big(\bm{\theta}^{k+1},\,\bm{x}^{k+1}\big)\in\partial G(\bm{y}^{k+1},\,\bm{\xi}^{k+1})$, we can get the following equality by using the Fenchel equality (see, for example, [35, Theorem 23.5]):
@@
Then, by passing to a further subsequence if necessary, we may also assume without loss of generality that the subsequence $\{\bm{x}^{k_{j}}\}\subseteq\{\bm{x}^{k}\}$ converges to some accumulation point $\bm{x}^{\infty}$, which, in view of statement (iv), belongs to $\mathcal{X}^{*}$ (the optimal solution set of (3.2)). Thus, for such $\bm{x}^{\infty}$ and any $\bm{y}^{*}\in\mathcal{Y}^{*}$, using (A.10) with some manipulations, we obtain, for all $k>k_{j}$,
@@
Since $0\leq\phi\leq\liminf\limits_{k\to\infty}\left\{\|\bm{w}^{k}-\bm{y}^{*}\|^{2}+\tau_{k}\|\bm{y}^{k}-\bm{y}^{*}\|^{2}\right\}$ for any $\bm{y}^{*}\in\mathcal{Y}^{*}$, passing to the limit superior when $k\to\infty$ on the both sides of this inequality, we obtain that, for any $\bm{y}^{*}\in\mathcal{Y}^{*}$,
@@
where $\mu_{k}:=\sqrt{\frac{1+\nu_{k}}{1+\gamma_{k}}}$. Since $\nu_{k}\to 0$ and $\gamma_{k}\geq\gamma_{\min}>0$ for all $k\color[rgb]{0,0,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.9}\geq k_{0}$, one can verify that $\limsup\limits_{k\to\infty}\,\{\mu_{k}\}<1$. Thus, we obtain the desired results in statement (i).
@@
$\widetilde{c}\,\tau_{\min}\|\bm{y}^{k+1}-{\bm{y}}^{k}\|^{2}\leq\widetilde{c}\left\|\begin{matrix}\sqrt{\tau_{k}}(\bm{y}^{k+1}-{\bm{y}}^{k})\\ \bm{x}^{k+1}-{\bm{x}}^{k}\end{matrix}\right\|^{2}\leq\mathrm{dist}_{\Lambda^{k}}^{\color[rgb]{0,0,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.9}2}\left((\bm{y}^{k},\bm{x}^{k}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right),$